OpenStudy (anonymous):

what are the zeros of 16x^4 -1

jimthompson5910 (jim_thompson5910):

Use the difference of squares rule to factor the left side \[\Large 16x^4 -1 = 0\] \[\Large (4x^2)^2 - (1)^2 = 0\] \[\Large (4x^2-1)(4x^2+1) = 0\] \[\Large ((2x)^2-(1)^2)(4x^2+1) = 0\] \[\Large (2x-1)(2x+1)(4x^2+1) = 0\] I'll let you finish up.

OpenStudy (anonymous):

still don't get it, thanks though

jimthompson5910 (jim_thompson5910):

you then set each factor equal to 0

jimthompson5910 (jim_thompson5910):

so you'll have these equations 2x-1 = 0 2x+1 = 0 4x^2+1 = 0 solve each for x